OpenStudy (megannicole51):

split into partial fractions....((5)/(y^3-25y))

zepdrix (zepdrix):

\[\Large \frac{5}{y^3-25y} \quad=\quad \frac{5}{y(y^2-5^2)}\]This might be a good first step right? :) Looks like we can further factor though. Remember how to break down the `difference of squares` ?

OpenStudy (megannicole51):

(A/y)+(B/y-5)+(C/(y-5)^2)?

zepdrix (zepdrix):

Woops you may have factored your denominator incorrectly, let's see.

zepdrix (zepdrix):

\[\Large \frac{5}{y(y^2-5^2)} \ne\frac{5}{y(y-5)^2}\]

OpenStudy (megannicole51):

thats how my professor broke down a problem in my notes....probably applies to a different type of problem

zepdrix (zepdrix):

yah probably D:

zepdrix (zepdrix):

We use this rule for factoring the difference of squares: \[\large \color{royalblue}{a^2-b^2=(a-b)(a+b)}\] So what does that do to our\[\Large y^2-25^2\]

zepdrix (zepdrix):

woops* \(\Large y^2-5^2\)

OpenStudy (megannicole51):

im confused....why are you doing that?

OpenStudy (anonymous):

watch the vid if you dont get it by the first vid there are 2 more vids :)

zepdrix (zepdrix):

We would like to be able to write the denominator as `linear factors` if we're able to. Which means all factors would contain y to the first power. We're starting with y^3 and trying to break it down to y^1's. It will make our partial fraction decomposition much easier.

OpenStudy (megannicole51):

@pgpilot326

hero (hero):

The first step is to factor the denominator: \[y^3-25y = y(y^2 - 25) = y(y + 5)(y - 5)\] Afterwards, you write the partial expansion in the form: \[\frac{5}{y(y + 5)(y - 5)} = \frac{A}{y - 5} + \frac{B}{y} + \frac{C}{y + 5}\]

OpenStudy (megannicole51):

i put that earlier and someone said it was wrong?

OpenStudy (megannicole51):

actually my numerators were a bit different

OpenStudy (anonymous):

look at the denominator and factor, right? \[ y^{3} -25y = y\left(y^{2}-25\right) = y(y-5)(y+5)\] now look over those papers and figure out what to do. if you get stuck, message me or reply addressing me. You've got this, girl!

hero (hero):

Next multiply both sides by \(y(y + 5)(y - 5))\), then simplify: \[5 = y(y + 5)A + (y - 5)(y + 5)B + y(y - 5)C\]

OpenStudy (megannicole51):

why did you set the denominators up the way u did?

hero (hero):

@megannicole51, because that's the form you want when expanding to partial fractions.

hero (hero):

Now you have to expand the right side

hero (hero):

The right side expands to: \(y(y + 5)A + (y + 5)(y - 5)B + y(y - 5)C\) \(= (y^2 + 5y)A + (y^2 - 25)B + (y^2 - 5y)C\) \(=Ay^2 + A5y + By^2 - 25B + Cy^2 - C5y\) \(=Ay^2 + By^2 + Cy^2 + A5y - C5y - 25B\) \(=(A + B + C)y^2 + (A - C)5y - 25B\)

hero (hero):

\( = (A + B + C)y^2 + (5A - 5C)y - 25B\)

OpenStudy (megannicole51):

i have to go to class but ill be back to take a look at ur expansion!

hero (hero):

Basically you end up with this: \(5 = (A + B + C)y^2 + (5A - 5C)y - 25B\) And then you set up a system like so: \(0 = A + B + C\) \(0 = 5A - 5C\) \(5 = -25B\)

hero (hero):

Then after solving the system you end up with: \[A = \frac{1}{10}\]\[B = -\frac{1}{5}\]\[C = \frac{1}{10}\]

hero (hero):

And the final form you get should be: \[\frac{5}{y(y + 5)(y - 5)} = \frac{1}{10(y - 5)} -\frac{1}{5y} + \frac{1}{10(y + 5)}\]

OpenStudy (megannicole51):

0=A+B+C 0=5A−5C 5=−25B how do you know this?

OpenStudy (megannicole51):

@Hero

OpenStudy (anonymous):

3eq with 3 unknowns : start with 5=−25B divide both sides by -25 B = -1/5 now from the second you get A=C plug B=-1/5 and A=C into the first 0=A-1/5+A 2A = 1/5 A = 1/10 nnow since A=C C=1/10

OpenStudy (anonymous):

oh that is not what you asked.

OpenStudy (megannicole51):

my question is just how do u know what those equal? 0=A+B+C 0=5A−5C 5=−25B

OpenStudy (anonymous):

so you had: 5 = (A+B+C)y^2 + (5A-5C)y -25B

OpenStudy (anonymous):

as you can see at the left side you have only a number (independent of y)

OpenStudy (anonymous):

so both of the terms that multiply y and y^2 at the right side must be 0

OpenStudy (anonymous):

A+B+C = 0 and 5A-5C = 0

OpenStudy (anonymous):

now what is left is the number at the right side : -25B it must be equal to the number at the left : 5 -25B = 5

OpenStudy (megannicole51):

oooh okay

OpenStudy (megannicole51):

Thank you to everyone who helped!

hero (hero):

Sorry, I stepped away from my computer for a moment. @megannicole51, do you get everything else?

OpenStudy (megannicole51):

yup im good!!! thank you! i just like working through these problems with people until i feel like ive got it:)